Green Island can either produce 20 tons of fish or 40 tons of coconuts. Putting fish on the x-axis, draw the PPF of Green Island. Calculate the opportunity costs of both goods in this country. Blue Peninsula can either produced 10 tons of fish or 20 tons of coconuts. Putting fish on the x-axis, draw. the PPF of Blue Peninsula Calculate the opportunity costs of both goods in this country. Will trade between these two countries be beneficial? Why
Answer #1

(1)

The PPF is the locus of all such points that represents the combination of two or more goods that a country can produce using all its resources efficiently. The figure1 below gives the PPF for this country. The opportunity cost of the good is the cost of producing one good is the amount of other good(s) foregone.

In this case, the country can either produce 20 tons of fish or 40 tons of coconut. Then to produce 20 tons fish, it gives up 40 tons of coconut. Then the opportunity cost of 20 tons of fish is 40 tons of coconut. The opportunity cost of 1 ton of fish is (40/20)=2 ton of coconut. Again, in order to produce 40 tons coconut, it gives up 20 tons of fish. Then the opportunity cost of 40 tons of coconut is 20 tons of fish. The opportunity cost of 1 ton of coconut is (20/40)=1/2 ton of fish.

PPF of Island 45 40 35 漠) 3 25 15 10 10 15 20 25 Fish

===================================================================================

(2)

The PPF is the locus of all such points that represents the combination of two or more goods that a country can produce using all its resources efficiently. The figure2 below gives the PPF for this country. The opportunity cost of the good is the cost of producing one good is the amount of other good(s) foregone.

In this case, the country can either produce 10 tons of fish or 20 tons of coconut. Then to produce 10 tons fish, it gives up 20 tons of coconut. Then the opportunity cost of 10 tons of fish is 20 tons of coconut. The opportunity cost of 1 ton of fish is (20/10)=2 ton of coconut. Again, to produce 20 tons coconut, it gives up 10 tons of fish. Then the opportunity cost of 20 tons of coconut is 10 tons of fish. The opportunity cost of 1 ton of coconut is (10/20)=1/2 ton of fish.

PPF of Penninsula 25 20 15 O 10 10 12 Fish

===================================================================================

(3)

The theory of comparative advantage states that the countries can gain from trade if they spacialize and trade. That is if a country can produce a good it has lower opportunity cost, it can gain from trading that good with the good it does not have advantage in.

In this case, however, both countries have the same opportunity cost of producing goods. Then they both cannot gain from trade according to comparative advantage.